GISMO run #7; 2nd pool run (November 2012)

Back to the GISMO main page


Schedule: 05.04. - 21.04.2012

Summarized schedule of the run:


Main event

Friday 26/10 - Tuesday 30/10

Preparation: cool down, system checks, calibration

Tuesday 30/10 - Tuesday 13/11

Calibration (1st day); observing pool; Warm up & switch off (last day)

GISMO final version

Daily reports

Friday 26/10

We have the electronics and computers setup and GISMO is on the pump. IRAM left the electronics in the cabin which gave us a head start. We also decided to put the cabling above the floor (with a good metal cover) to reduce the chances of breaking thermometry cables while fishing cables and also to make setup/teardown easier. We also cleaned up the IRIG isolation cabling and put the isolator box in the rack. The server was a little stubborn to boot but eventually gave in. The server, client, crate, IRIG, are up and running. We also made some fresh clones of the server disk and tested the backup server (booted and started GISMO software).

Saturday 27/10

Continuing without problem. Weather getting a little better. Continuing cooldown. Added nitrogen. Prepared GISMO Nexus for new run. Removed ground breaker from electronics circuit. Completed disk backups and setting up client. Elmer working on optical models and helium in-gassing prevention design.

Monday 29/10

Completed cooldown (took lots of IV curves during extension cooldown). Telecon with IRAM. Added more thermometry (extension) to Nexus/alarm system. Took some initial observations during technical time. We had a problem with IDL license (the error was the license file does not exist - as opposed to last spring where there were too many users). We copied the license file from gra-lx1 and that worked. Also, CRUSH now(?) wants headers in initial log file, so we obliged. Finally, we ran into gismo directory permission problems and emailed Walter. After clearing things up we manually sent the data through the backend - see [[GISMO Nexus|])

Thursday 01/11

Yesterday we made a few tests during the bad weather with the window blanked off. The aluminum plate was bolted to the bottom of the cryostat and we configured GISMO with different types of grounding schemes to find the lowest possible noise. After collecting many data sets, we discovered that the best configuration is with all of the electronics connected to the electrical ground in the cabin and the cryostat vacuum shell connected to isolated ground. The isolated ground is a special cable feed into the receiver cabin and has the closest possible contact to the earth ground for the telescope. In this configuration, the noise was lowered.

Since the bad weather is continuing today, we will try to make further grounding tests.

Friday 02/11 - Saturday 03/11

We observed about 15hours so far, being hindered by a lot of poor weather. The forecast is bad for the coming days. From the GISMO team Elmer and Steve did leave on Thursday and Friday, as planned, while Attila and Johannes arrived.

We encounter some problems with the cryostat. The NEFDs have been a bit higher as during the last run, 15-12 mJy sqrt(s). In addition, from the behavior of the vacuum (and temperatures), Santiago and Johannes conclude that there should be a leak somewhere with Helium freezing out inside. Juan Luis and Santiago did a leak test this morning, and identified a leak near the window. During the last two days, we warmed-up slightly twice to get rid of the Helium.

Otherwise, pointing and focus are fine. Observations run smoothly when the weather allows.

Sunday 04/11

We failed to fix the vacuum leak at the window seal with vacuum grease and tape. As all weather forecasts continue to be very bad till Friday, we decided right now to warm-up and clean the window or even replace it. We have material for windows here, Santiago will stay on the mountain, and Elmer is helping us via skype. By Monday evening the dewar should be warm and about 30hours later we should be operational again, i.e. on Wednesday. The advantage of working on the window now, is that we will be able to measure sensitivity later during the week, to confirm that GISMO is fine. Any new window will have the same 3/4 wavelength thickness and will be the same material (pp) as the current one.

Right now, we have strong winds of upto 41m/s and heavy rain.

Monday 05/11 - Tuesday 06/11

We dismounted the window, found some cracks, build a new window, put it in place and pumped on the cryostat.

Next morning, the cryostat being under vacuum we proceed to a new leak test. Unfortunatelly there's still a leak, apparently near the seal of the plate holding the window; so a bit higher than we previouly thought. We could live with that for now, or since the weather will still be bad for the comming days we could also decide to open further the cryostat to fix this leak.

Removed the mirror, opened the bottom plate of the baffle, cleaned o-ring and connecting surfaces, closed baffle again, pumped down, no more leaks found, cooled down.

Wednesday 07/11 - Thursday 08/11

Remove nitrogen from the He tank and fill in liquid He at noon. At the end of the afternoon the detectors were cold and ready to observe, but the wind didn't allow it. An intervention and reboot of the GISMO server has been necessary caused by a double login. Blanked off and then opened to the cabin we measured an unxecped low flux of photons on the detectors, this was caused by the unxepected presence of one of the movable filter into the beam; once put out, the behaviour of the TES looked exactly as expected. GISMO is in perfect shape.

Next morning again a need to reboot the GISMO server. Since the weather still not alowed us to do observations we tried to create fake observations using the Cal command in order to generate data that can be processed by CRUSH and give the full picture of the performances. Unfortunatelly it has not been anticipated that the Cal command would be used with Pako in the GISMO configuration, and when doing so NCS generated incomplete IMBFITs unusable by crush; only an IRAM expert can fix this.

Friday 09/11

Strong Wind. Clouds, rain, snow. :-(

After operators switched to demo mode, and after work by WB, we were able to run normal lissajous scripts while the telescope stayed in stow position. These scans were processed by the pipeline, merged fits files were created. However, crush complaints and delivers weird NEFDs in the nanoJy range with GISMO closed by the cap. Crush must be confused by one of the parameters, even though we used the vclip option.

Saturday 10/11

At last the sky cleared up sufficiently during the night to be able to finally resume observations ! In the morning we had to stop again due to clouds, but since mid morning we are observing, with some interruptions due to some clouds passing. On the clear sky when observing the Galactic Center the performances of GISMO are excellent as expected, with sensitivities around 10mJy*s^0.5. Other observations showed about 15mJy (always referring to the results of the default crush pipeline creating the log).

Our operation on fixing the leak was successful; the He consumption of the cryostat has significantly decreased and we can now run 24h before having to recyle the absorption coolers and to refill the He and N2 tanks. Note: the leak had an effect only on the He consumption, not on the detector performances.

Sunday 11/11


Monday 12/11

At last we finally had some decent weather allowing observations in the afternoon. And everything went fine during these few hours. Wind too strong in the evening, observations stopped. 3h latter: significant wind decreased, clear sky, resume observations.

GismoRun7 (last edited 2012-11-13 00:39:58 by NikaBolometer)